If an object M1=40kg elastically collides with another object m2=30kg which happens to be initially at rest, what would be the velocity of m1 before and after the collision.

9l60 Pots y the CollsiOn 0 2 ,略 offe)
Answer #1

given data

m1 = 40 kg

m2 = 30 kg

v2 = 10 m/s

ui=?

v1 =?

it is an elastic collision

so, From teh conservation of linear momentum

Pi = Pf

resolve the velocity components

horizontally Pi = Pf

m1*u1 cos(30)+0 = m1*v1*cos(60)- m2*v2 *sin(25)

40*u1*0.866 = 40*v1 *0.5 - 30*10*0.422

34.64 u1 = 20*v1 - 126.6-----------------------(1)

vertically,

m1*u1 sin(30)+0 = m1*v1*sin(60)- m2*v2 *cos(25)

40*u1*0.5 + 0= 40*v1*0.866 - 30 * 10*0.906

20 u1 = 34.64*v1 - 271.8------------(2)

From (1)

u1 = (20*v1 - 126.6)/(34.64) substitute in (2)

20*( (20*v1 - 126.6)/(34.64)) = 34.64*v1 - 271.8

v1 = 8.604 m/s after collision

so,, 20 u1 = 34.64*8.604 - 271.8

u1 = 1.312 m/s Before collision.